OpenStudy (anonymous):

The most easiest question ever. If you can do it correctly, I will give you a medal. What's 2+2=?

OpenStudy (anonymous):

lol its 4

OpenStudy (austinl):

2+2=5 2+2=2+2+0 x = y x^2 = xy x^2-y^2 = xy-y^2 (x+y)(x-y) = y(x-y) x+y = y 2y = y 2 = 1 1 = 0 Therefore, 2+2+0=2+2+1=5

OpenStudy (dobby1):

4

OpenStudy (anonymous):

on solving this as per the rules of algebraic expressions, we get x/y=0 and x=y is not suitable

OpenStudy (austinl):

Robins Lemma proof.

OpenStudy (austinl):

But yes, there is some "fuzzy math" there.

OpenStudy (anonymous):

but this is against assumptions

OpenStudy (anonymous):

Pretty sure it's FISH 2+2=FISH

OpenStudy (anonymous):

The winner is...Blakarican. S/he showed me how to do it, and gave me the CORRECT answer. This...was a trick question.

OpenStudy (anonymous):

Haha @austinL Im boss lol

OpenStudy (anonymous):

Thankyou sir for the medal :D

OpenStudy (anonymous):

You're welcome!

OpenStudy (austinl):

@giligails If we have, \(x=y\) \(\dfrac{x}{y}=1\) By the way...

OpenStudy (anonymous):

@Shawna.marie @Asuna_Yuuki

OpenStudy (anonymous):

Guess what it 2+2=sandwich

OpenStudy (anonymous):

2+2=fish lol fairly odd parents #Ihavenolife

OpenStudy (anonymous):

Ok, thank you everyone! No more answers are required!